# Adaptive-TMLE for the Average Treatment Effect based on Randomized Controlled Trial Augmented with Real-World Data

### v1

@@ 7 Asymptotic Super-Efficiency of A-TMLE
Moreover, under the same Donsker class condition, using the consistency of $P_{n}^{*}$, it also follows that
$\Psi_{{\cal M}_{w,n}}(P_{n}^{*})-\Psi(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w,n},P_{0}}+o_{P}(n^{-1/2}).$
@@
Finally, if $D_{{\cal M}_{w,n},P_{0}}$ remains random in the limit but is asymptotically independent of the data $P_{n}$ so that $n^{1/2}P_{n}D_{{\cal M}_{w,n},P_{0}}/\sigma_{0,n}\Rightarrow_{d}N(0,1)$ with $\sigma_{0,n}^{2}\equiv P_{0}D_{{\cal M}_{w,n},P_{0}}^{2}$, then we obtain
$\sigma_{0,n}^{-1}n^{1/2}\left(\Psi_{{\cal M}_{w,n}}(P_{n}^{*})-\Psi(P_{0})\right)\Rightarrow_{d}N(0,1),$
@@
thereby allowing for construction of confidence intervals for $\Psi(P_{0})$. If, in fact, $D_{{\cal M}_{w,n},P_{0}}$ converges to a fixed $D_{{\cal M}_{0},P_{0}}$, then we have asymptotic linearity:
$\Psi_{{\cal M}_{w,n}}(P_{n}^{*})-\Psi(P_{0})=P_{n}D_{{\cal M}_{0},P_{0}}+o_{P}(n^{-1/2}),$
with influence curve the efficient influence curve of $\Psi_{{\cal M}_{0}}:{\cal M}\rightarrow\hbox{${\rm I\kern-1.99997ptR}$}$ at $P_{0}$. Moreover, [adml] shows that $D_{{\cal M}_{0},P_{0}}$ equals the efficient influence curve of $\Psi:{\cal M}_{0}\rightarrow\hbox{${\rm I\kern-1.99997ptR}$}$ that a priori assumes the model ${\cal M}_{0}$, showing that this adaptive-TMLE is super-efficient achieving the efficiency bound for estimation $\Psi(P_{0})$ under the oracle model ${\cal M}_{0}$ (as if we are given $P_{0}\in{\cal M}_{0}$ for given ${\cal M}_{0}$).
@@
Thus, if the true $\tau_{P_{0}}$ is captured by a small model, then the adaptive-TMLE will achieve a large gain in efficiency relative to the regular TMLE, while if $\tau_{P_{0}}$ is complex so that ${\cal M}_{0}$ is close to ${\cal M}$, then the gain in efficiency will be small or the adaptive-TMLE will just be asymptotically efficient. The discussions above are summarized into the following theorem.
@@
Let $D_{{\cal M}_{w,n},P}$ be the canonical gradient of $\Psi_{{\cal M}_{w,n}}=\Psi(\Pi(\cdot\mid{\cal M}_{w,n}))$ at $P$. Let $R_{{\cal M}_{w,n}}(P,P_{0})\equiv\Psi_{{\cal M}_{w,n}}(P)-\Psi_{{\cal M}_{w,n}}(P_{0})+P_{0}D_{{\cal M}_{w,n},P}$ be the exact remainder. Let ${\cal M}_{0}\subset{\cal M}$ contain $P_{0}$, a so-called oracle model, and $D_{{\cal M}_{0},P}$ and $R_{{\cal M}_{0}}(P,P_{0})$ be the canonical gradient and exact remainder of $\Psi_{{\cal M}_{0}}:{\cal M}\rightarrow\hbox{${\rm I\kern-1.99997ptR}$}$, respectively. Let $P_{n}^{*}\in{\cal M}_{w,n}$ be an estimator of $P_{0}$ such that $P_{n}D_{{\cal M}_{w,n},P_{n}^{*}}=o_{P}(n^{-1/2})$. Then,
$\Psi_{{\cal M}_{w,n}}(P_{n}^{*})-\Psi_{{\cal M}_{w,n}}(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w,n},P_{n}^{*}}+R_{{\cal M}_{w,n}}(P_{n}^{*},P_{0}).$
Oracle model approximation condition: Let $P_{0,n}\equiv\Pi_{{\cal M}_{w,n}}(P_{0})$ and $\tilde{P}_{0,n}=\Pi(P_{0,n}\mid{\cal M}_{0})$. Let $\tilde{D}_{{\cal M}_{w,n},P_{0,n}}$ be the projection of $D_{{\cal M}_{0},P_{0,n}}$ onto the tangent space $T_{{\cal M}_{w,n}}(P_{0,n})\subset L^{2}_{0}(P_{0,n})$ of the working model ${\cal M}_{w,n}$ at $P_{0,n}$. Assume that ${\cal M}_{w,n}$ approximates ${\cal M}_{0}$ in the sense that $\Psi(\tilde{P}_{0,n})-\Psi(P_{0,n})=o_{P}(n^{-1/2})$ and
$(\tilde{P}_{0,n}-P_{0})\left\{D_{{\cal M}_{0},\tilde{P}_{0,n}}-\tilde{D}_{{\cal M}_{w,n},\tilde{P}_{0,n}}\right\}=o_{P}(n^{-1/2}).$
@@
$\Psi_{{\cal M}_{w,n}}(P_{n}^{*})-\Psi(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w,n},P_{n}^{*}}+R_{{\cal M}_{w,n}}(P_{n}^{*},P_{0})+o_{P}(n^{-1/2}).$
Rate of convergence condition: Under the condition that $P_{n}^{*}$ converges to $P_{0}$ at a fast enough rate (i.e., $o_{P}(n^{-1/4})$ so that $R_{{\cal M}_{w,n}}(P_{n}^{*},P_{0})=o_{P}(n^{-1/2})$, this yields
$\Psi_{{\cal M}_{w,n}}(P_{n}^{*})-\Psi(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w,n},P_{n}^{*}}+o_{P}(n^{-1/2}).$
Donsker class condition: Under the Donsker class condition that $D_{{\cal M}_{w,n},P_{n}^{*}}$ falls in a $P_{0}$-Donsker class with probability tending to 1, this implies $\Psi_{{\cal M}_{w,n}}(P_{n}^{*})-\Psi(P_{0})=O_{P}(n^{-1/2})$, consistency at the parametric rate $n^{-1/2}$. Moreover, under the same Donsker class condition, using the consistency $P_{0}\{D_{{\cal M}_{w,n},P_{n}^{*}}-D_{{\cal M}_{w,n},P_{0}}\}^{2}\rightarrow_{p}0$ of $P_{n}^{*}$, it also follows that
$\Psi_{{\cal M}_{w,n}}(P_{n}^{*})-\Psi(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w,n},P_{0}}+o_{P}(n^{-1/2}).$
Asymptotic normality condition: Finally, if $D_{{\cal M}_{w,n},P_{0}}$ remains random in the limit but is asymptotically independent of the data $P_{n}$ in the sense that $n^{1/2}P_{n}D_{{\cal M}_{w,n},P_{0}}/\sigma_{0,n}\Rightarrow_{d}N(0,1)$ with $\sigma_{0,n}^{2}\equiv P_{0}D_{{\cal M}_{w,n},P_{0}}^{2}$, then we obtain
$\sigma_{0,n}^{-1}n^{1/2}\left(\Psi_{{\cal M}_{w,n}}(P_{n}^{*})-\Psi(P_{0})\right)\Rightarrow_{d}N(0,1),$
thereby allowing for construction of confidence intervals for $\Psi(P_{0})$. If, in fact, $D_{{\cal M}_{w,n},P_{0}}$ converges to a fixed $D_{{\cal M}_{0},P_{0}}$, then we have asymptotic linearity:
$\Psi_{{\cal M}_{w,n}}(P_{n}^{*})-\Psi(P_{0})=P_{n}D_{{\cal M}_{0},P_{0}}+o_{P}(n^{-1/2}),$
@@ 8 Simulation Studies
We conducted two sets of simulations to evaluate our A-TMLE estimator against alternative methods. The first simulation setting involves augmenting both the treatment and control arm. The second simulation is a scenario with rare binary outcomes using only control data from the real-world, a situation likely when treatment data are unavailable, for example, due to pending drug approvals and the rarity of the outcome could lead to an underpowered RCT. Additional simulations for more general data structures with missing outcomes are available in Appendix B. Descriptions of the five estimators we assessed is shown in Table 1.
@@
Our primary metrics for evaluating the performance of the estimators include mean-squared-error (MSE), relative MSE, coverage and width of 95% confidence intervals.
8.1 Augmenting both the treatment and control arm
@@
Scenarios (a) and (b) have bias functions (specifically, the conditional effect of the study indicator on the outcome) of simple main-term linear forms, while in scenarios (c) and (d) we introduced complex biases involving indicator jumps, higher-order polynomial terms and interactions, demonstrating the misspecification of a main-term linear model and the necessity of employing HAL as a flexible nonparametric regression algorithm for accurate approximation of the bias function. In scenarios (a) and (b), we constructed a hybrid design with the external data sample size being threefold that of the RCT, mirroring the common scenario where, for example, electronic health records databases offer a substantially larger pool of external data. The external dataset comprises both treated and control patients. In the RCT, the probability of assignment to the active treatment group is 0.67, reflecting realistic scenarios where a drug has been approved and the focus is on assessing safety for secondary outcomes. Treatment assignment in the external data mimics real-world conditions, determined by baseline patient characteristics, as would be the case in clinical practice. To ensure that a straightforward pooling of the two data sources would yield a biased estimate of the true ATE, we added a bias term to the outcome of the patients from the external data.
@@
Figure 2: Comparing the MSE, relative MSE, 95% confidence interval coverage of A-TMLE, ES-CVTMLE, and TMLE estimator for increasing sample sizes of both the RCT and external data. The reference (red dashed line) in the relative MSE plots are the MSEs of the RCT-only estimator. The plots in the right-most column show the average proportion of cross-validation folds the ES-CVTMLE estimator pooled the external data.
@@
Figure 2 demonstrates that the A-TMLE estimator (red line) has significantly lower MSE than both ES-CVTMLE and TMLE across scenarios (a) and (b), being 1.5 times more efficient than an efficient estimator that uses RCT data alone. All estimators maintain nominal 95% confidence interval coverage. Notice that, in scenario (a), the smaller bias from external data allows ES-CVTMLE (green line) to achieve moderate efficiency gains over TMLE (blue line), with the pooled parameter selected in about 70% of its cross-validation folds, indicating its ability to utilize external data for efficiency improvements. Despite this, A-TMLE surpasses ES-CVTMLE in efficiency gains. Scenario (b) amplifies the bias magnitude introduced from the external data, severely diminishing ES-CVTMLE’s efficiency gain, nearly to nonexistence asymptotically. Conversely, A-TMLE’s capability to accurately learn the bias working model data-adaptively enhances its efficiency significantly, even as the sample size increases and despite the larger bias, maintaining robust type 1 error control. ES-CVTMLE’s failure to achieve efficiency gain in scenario (b) is highlighted by its near-total rejection of the external data due to large estimated bias. Recall that in the Introduction section, we mentioned that an efficient estimator for the target parameter $\Psi$ may still offer no gain in the presence of external data, this point is demonstrated in the relative MSE plots with an efficient TMLE that goes after $\Psi$ (blue line) matching the MSE of an estimator using RCT data only (red dashed line). Therefore, even theoretically optimal estimators may not yield efficiency gains from external data integration.
In some practical settings, bias may deviate from the simple main-term linear model we considered in our first simulation, manifesting as a complex function of baseline covariates and treatment. This complexity necessitates the employment of more flexible nonparametric learning algorithms for accurate bias estimation. Hence, our second simulation utilizes HAL with a rich set of spline basis functions to data-adaptively learn the bias working model. Echoing the design of the first simulation, we increase the external data’s sample size to five times that of the RCT, maintaining a 0.67 treatment randomization probability within the RCT, and both treated and control patients are present in the external data. The bias term in this scenario includes indicator jumps, higher-order polynomials, and interaction terms, rendering a simple main-term linear model misspecified. The data-generating process is provided in Appendix B.

### v2

@@ 6 Asymptotic Super-Efficiency of A-TMLE
Moreover, under the same Donsker class condition, using the consistency of $P_{n}^{\star}$, it also follows that
$\Psi_{{\cal M}_{w}}(P_{n}^{\star})-\Psi(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w},P_{0}}+o_{P}(n^{-1/2}).$
@@
Finally, if $D_{{\cal M}_{w},P_{0}}$ remains random in the limit but is asymptotically independent of the data $P_{n}$ so that $n^{1/2}P_{n}D_{{\cal M}_{w},P_{0}}/\sigma_{0,n}\Rightarrow_{d}N(0,1)$ with $\sigma_{0,n}^{2}\equiv P_{0}D_{{\cal M}_{w},P_{0}}^{2}$, then we obtain
$\sigma_{0,n}^{-1}n^{1/2}\left(\Psi_{{\cal M}_{w}}(P_{n}^{\star})-\Psi(P_{0})\right)\Rightarrow_{d}N(0,1),$
@@
thereby allowing for the construction of Wald-type confidence intervals for $\Psi(P_{0})$. If, in fact, $D_{{\cal M}_{w},P_{0}}$ converges to a fixed $D_{{\cal M}_{0},P_{0}}$, then we have asymptotic linearity:
$\Psi_{{\cal M}_{w}}(P_{n}^{\star})-\Psi(P_{0})=P_{n}D_{{\cal M}_{0},P_{0}}+o_{P}(n^{-1/2}),$
@@
with influence curve the efficient influence curve of $\Psi_{{\cal M}_{0}}:{\cal M}\rightarrow\hbox{${\rm I\kern-1.99997ptR}$}$ at $P_{0}$. Moreover, [adml] shows that $D_{{\cal M}_{0},P_{0}}$ equals the efficient influence curve of $\Psi:{\cal M}_{0}\rightarrow\hbox{${\rm I\kern-1.99997ptR}$}$ that a priori assumes the model ${\cal M}_{0}$, showing that this A-TMLE is super-efficient achieving the efficiency bound for estimation $\Psi(P_{0})$ under the oracle model ${\cal M}_{0}$ (as if we are given $P_{0}\in{\cal M}_{0}$ for given ${\cal M}_{0}$). Thus, if the true $\tau_{A,0}$ and $\tau_{S,0}$ are captured by simple models, then the A-TMLE will achieve a large gain in efficiency relative to the regular TMLE, while if $\tau_{A,0}$ and $\tau_{S,0}$ complex so that ${\cal M}_{0}$ is close to ${\cal M}$, then the gain in efficiency will be small or the resulting A-TMLE will just be asymptotically efficient. The discussions above are summarized into the following theorem.
@@
Let $D_{{\cal M}_{w},P}$ be the canonical gradient of $\Psi_{{\cal M}_{w}}=\Psi(\Pi_{{\cal M}_{w}}(\cdot))$ at $P$. Define the exact remainder as $R_{{\cal M}_{w}}(P,P_{0})\equiv\Psi_{{\cal M}_{w}}(P)-\Psi_{{\cal M}_{w}}(P_{0})+P_{0}D_{{\cal M}_{w},P}$. Let ${\cal M}_{0}\subset{\cal M}$ denote the oracle model containing $P_{0}$, with $D_{{\cal M}_{0},P}$ and $R_{{\cal M}_{0}}(P,P_{0})$ being the canonical gradient and exact remainder of $\Psi_{{\cal M}_{0}}:{\cal M}\rightarrow\hbox{${\rm I\kern-1.99997ptR}$}$, respectively. Let $P_{n}^{\star}\in{\cal M}_{w}$ be an estimator of $P_{0}$ such that $P_{n}D_{{\cal M}_{w},P_{n}^{\star}}=o_{P}(n^{-1/2})$. Then,
$\Psi_{{\cal M}_{w}}(P_{n}^{\star})-\Psi_{{\cal M}_{w}}(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w},P_{n}^{\star}}+R_{{\cal M}_{w}}(P_{n}^{\star},P_{0}).$
Oracle model approximation condition: Let $P_{0,n}\equiv\Pi_{{\cal M}_{w}}(P_{0})$ and $\tilde{P}_{0,n}\equiv\Pi_{{\cal M}_{0}}(P_{0,n})$. Let $\tilde{D}_{{\cal M}_{w},P_{0,n}}$ be the projection of $D_{{\cal M}_{0},P_{0,n}}$ onto the tangent space $T_{{\cal M}_{w}}(P_{0,n})\subset L^{2}_{0}(P_{0,n})$ of the working model ${\cal M}_{w}$ at $P_{0,n}$. Assume that ${\cal M}_{w}$ approximates ${\cal M}_{0}$ in the sense that $\Psi(\tilde{P}_{0,n})-\Psi(P_{0,n})=o_{P}(n^{-1/2})$ and
$(\tilde{P}_{0,n}-P_{0})\left\{D_{{\cal M}_{0},\tilde{P}_{0,n}}-\tilde{D}_{{\cal M}_{w},\tilde{P}_{0,n}}\right\}=o_{P}(n^{-1/2}).$
@@
$\Psi_{{\cal M}_{w}}(P_{n}^{\star})-\Psi(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w},P_{n}^{\star}}+R_{{\cal M}_{w}}(P_{n}^{\star},P_{0})+o_{P}(n^{-1/2}).$
Rate of convergence condition: If $P_{n}^{\star}$ converges to $P_{0}$ at a fast enough rate (i.e., $o_{P}(n^{-1/4})$ so that $R_{{\cal M}_{w}}(P_{n}^{\star},P_{0})=o_{P}(n^{-1/2})$, then we have
$\Psi_{{\cal M}_{w}}(P_{n}^{\star})-\Psi(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w},P_{n}^{\star}}+o_{P}(n^{-1/2}).$
Donsker class condition: Under the Donsker class condition that $D_{{\cal M}_{w},P_{n}^{\star}}$ falls in a $P_{0}$-Donsker class with probability tending to 1, it follows that $\Psi_{{\cal M}_{w}}(P_{n}^{\star})-\Psi(P_{0})=o_{P}(n^{-1/2})$, consistency at the parametric rate $n^{-1/2}$. Moreover, under the same Donsker class condition, using the consistency $P_{0}\{D_{{\cal M}_{w},P_{n}^{\star}}-D_{{\cal M}_{w},P_{0}}\}^{2}\rightarrow_{p}0$ of $P_{n}^{\star}$, it also follows that
$\Psi_{{\cal M}_{w}}(P_{n}^{\star})-\Psi(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w},P_{0}}+o_{P}(n^{-1/2}).$
Asymptotic normality condition: Finally, if $D_{{\cal M}_{w},P_{0}}$ remains random in the limit but is asymptotically independent of the data $P_{n}$ in the sense that $n^{1/2}P_{n}D_{{\cal M}_{w},P_{0}}/\sigma_{0,n}\Rightarrow_{d}N(0,1)$ with $\sigma_{0,n}^{2}\equiv P_{0}D_{{\cal M}_{w},P_{0}}^{2}$, then we obtain
$\sigma_{0,n}^{-1}n^{1/2}\left(\Psi_{{\cal M}_{w}}(P_{n}^{\star})-\Psi(P_{0})\right)\Rightarrow_{d}N(0,1),$
thereby allowing for the construction of Wald-type confidence intervals for $\Psi(P_{0})$. If, in fact, $D_{{\cal M}_{w},P_{0}}$ converges to a fixed $D_{{\cal M}_{0},P_{0}}$, then we have asymptotic linearity:
$\Psi_{{\cal M}_{w}}(P_{n}^{\star})-\Psi(P_{0})=P_{n}D_{{\cal M}_{0},P_{0}}+o_{P}(n^{-1/2}),$
@@
7 Implementation of A-TMLE
In this section, we describe the implementation of A-TMLE for data augmentation and discuss practical strategies for data-adaptively learning the working models. Broadly, the objective is to construct two estimators, one for the pooled-ATE estimand and the other for the bias estimand. The pooled-ATE is estimated by combining the RCT data with the external RWD. However, direct pooling is likely to introduce bias, requiring a subsequent step to estimate and correct this bias. The bias can be expressed as a weighted combination of the conditional average RCT-enrollment effects for the treatment and control arms. Estimating the bias estimand parallels the estimation of an ATE, with trial enrollment acting as the treatment variable. The final estimate of the target parameter is then obtained by subtracting the bias estimate from the pooled-ATE estimate. The complete procedure for implementing the A-TMLE algorithm is detailed in Algorithm 1.
Algorithm 1 Adaptive-TMLE
1:Construct initial estimator $\tilde{\Psi}_{\mathcal{M}_{A,w}}(P_{n})$ of $\tilde{\Psi}_{\mathcal{M}_{A,w}}(P_{0})$:
2:     Compute nuisance estimate $g_{n}$ of $g_{0}(1\mid W)\equiv P_{0}(A=1\mid W)$;
3:     Compute nuisance estimate $\theta_{n}$ of $\theta_{0}(W)\equiv E(Y\mid W)$;
4:     Define pseudo outcome $\tilde{Y}\equiv(Y-\theta_{n})/(A-g_{n})$ and weight $\tilde{\mathcal{W}}\equiv(A-g_{n})^{2}$;
5:     Fit a relaxed-HAL using $W$ as covariates, $\tilde{Y}$ as outcome, and $\tilde{\mathcal{W}}$ as weight;
6:     Get initial estimate $\tilde{\Psi}_{\mathcal{M}_{A,w}}(P_{n})=1/n\sum_{i=1}^{n}\tau_{A,\beta_{n}^{\star}}(W_{i})=\tilde{\Psi}_{\mathcal{M}_{A,w}}(P_{n}^{\star})$;
7:Construct initial estimator $\Psi_{\mathcal{M}_{S,w}}^{\#}(P_{n})$ of $\Psi_{\mathcal{M}_{S,w}}^{\#}(P_{0})$:
8:     Obtain nuisance estimate $\Pi_{n}$ of $\Pi_{0}(1\mid W,A)\equiv P(S=1\mid W,A)$;
9:     Obtain nuisance estimate $\bar{Q}_{n}$ of $\bar{Q}_{0}(W,A)\equiv E(Y\mid W,A)$;
10:     Compute pseudo outcome $Y^{\#}\equiv(Y-\bar{Q}_{n})/(S-\Pi_{n})$ and weight $\mathcal{W}^{\#}\equiv(S-\Pi_{n})^{2}$;
11:     Fit a relaxed-HAL using $(W,A)$ as covariates, $Y^{\#}$ as outcome, and $\mathcal{W}^{\#}$ as weight;
12:     Get initial estimate $\Psi_{\mathcal{M}_{S,w}}^{\#}(P_{n})=1/n\sum_{i=1}^{n}[\Pi_{n}(0\mid W_{i},0)\tau_{S,\beta_{n}^{\star}}(W_{i},0)-\Pi_{n}(0\mid W_{i},1)\tau_{S,\beta_{n}^{\star}}(W_{i},1)]$;
13:Additional targeting for $\Pi$-component:
14:     Compute clever covariate $C(g_{n},\beta_{n}^{\star})=A/g_{n}(1\mid W)\tau_{S,\beta_{n}^{\star}}(W,1)-(1-A)/g_{n}(0\mid W)\tau_{S,\beta_{n}^{\star}}(W,0)$;
15:     Use $C(g_{n},\beta_{n}^{\star})$ to perform a TMLE targeting and obtain an updated estimate $\Pi^{\star}_{n}$ of $\Pi_{0}$;
16:     Get TMLE $\Psi_{\mathcal{M}_{S,w}}^{\#}(P_{n}^{\star})\equiv 1/n\sum_{i=1}^{n}[\Pi_{n}^{\star}(0\mid W_{i},0)\tau_{S,\beta_{n}^{\star}}(W_{i},0)-\Pi_{n}^{\star}(0\mid W_{i},1)\tau_{S,\beta_{n}^{\star}}(W_{i},1)]$;
17:Compute the final estimate $\Psi_{\mathcal{M}_{w}}(P_{n}^{\star})=\tilde{\Psi}_{\mathcal{M}_{A,w}}(P_{n}^{\star})-\Psi^{\#}_{{\cal M}_{S,w}}(P_{n}^{\star})$;
18:Compute the 95% confidence interval given by $\Psi_{\mathcal{M}_{w}}(P_{n}^{\star})\pm 1.96\sqrt{P_{n}(D_{\mathcal{M}_{A,w},\tilde{\Psi}}-D_{\mathcal{M}_{S,w},\Psi^{\#}})^{2}/n}$.
We make the following remarks regarding the algorithm. To estimate nuisance parameters, one might use either highly adaptive lasso (HAL) [benkeser_hal_2016, hejazi_hal_2020], or super learner [vdl_sl_2007]. Super learner, in particular, allows the integration of a rich library of flexible machine learning algorithms. For a practical guidance on specifying a super learner, we recommend [phillips_practical_2023]. For learning the two working models, we suggest using relaxed-HAL with the loss functions introduced in Sections 4 and 5. The pseudo-outcome and weight transformations defined in Algorithm 1 result from rewriting the loss function. Relaxed-HAL involves initially fitting a HAL where the $L_{1}$-norm of the coefficients is penalized. Cross-validation is used to select the optimal $L_{1}$-norm. After that, an ordinary least squares (OLS) regression is applied to the spline basis functions with non-zero coefficients. Because relaxed-HAL is a maximum likelihood estimator (MLE), the empirical mean of the $\beta$-component of the canonical gradient is automatically solved. This eliminates the need to further target the initial estimator $\tilde{\Psi}_{\mathcal{M}_{A,w}}$. Notably, HAL with 0th-order spline basis functions converges in loss-based dissimilarity at a rate of $n^{-1/3}(\log n)^{2(d-1)/3}$ [bibaut_fast_2019], only assuming that the target functions are càdlàg (right-continuous with left limits) and have finite sectional variation norm. Importantly, this rate depends on the dimensionality $d$ only via a power of the $\log n$-factor thus is essentially dimension-free. This allows the learned working model to approximate the truth well, enabling valid inference. Moreover, if the target functions are smooth, one can use $k$-th order spline HAL-MLEs that converge at even better rates $n^{-k^{\star}/(2k^{\star}+1)},k^{\star}=k+1$, up till a $\log n$-factor: for example, the first-order spline HAL-MLE converges at the dimension-free rate $n^{-2/5}$ up till a $\log n$-factor.
In Appendix D, we also discussed several alternative approaches for constructing data-adaptive working models, including learning the working model on independent data sets, and using deep learning architectures to generate dimension reductions of data sets. To complement the methods discussed in Appendix D, we also proposed and analyzed a cross-validated version of A-TMLE in Appendix C to address concerns regarding potential overfitting of the working models as a result of being too data-adaptive. In addition, as suggested in the Introduction, opting for a slightly larger model than indicated by cross-validation could reduce bias. Techniques like undersmoothing as described in [vdl_efficient_2023] are worth considering. Furthermore, to ensure robustness in model selection, one might enforce a minimal working model based on domain knowledge, specifying essential covariates that must be included in the working model. In the context of HAL, this would mean unpenalizing certain covariates or basis functions. This option is available in standard lasso software like the ‘glmnet’ R package [friedman_regularization_2010].
To estimate the nuisance parameters, one can use a super learner with a library of flexible machine learning algorithms. A discrete super learner, in particular, acts as an “oracle selector,” meaning that asymptotically it selects the algorithm that performs best within the library. This property holds even when the number of learners grows polynomially with the sample size, enabling the inclusion of a rich and diverse library of learners. If the super learner library includes HAL as a candidate learner, using it to estimate the nuisance parameters guarantees efficiency due to HAL’s fast rate of convergence. We note that proving efficiency also requires satisfying a Donsker class condition. This condition can often be bypassed by using cross-fitting. However, since HAL is an empirical risk minimizer within a cádlág function class, which is known to be Donsker, this condition is automatically satisfied when using HAL. In our simulations and real-data application, we applied cross-fitting when estimating the nuisance parameters. Intuitively, cross-fitting the nuisance parameters prevents overfitting, ensuring sufficient signal remains to accurately learn the working models for the conditional average treatment effect and the conditional average RCT-enrollment effect.
@@ 8 Simulation Studies
An ideal estimator in our problem setting should have several important operating characteristics including valid type I error control, nominal confidence interval coverage, and finite sample efficiency gain (compared to using RCT data only). We conducted two sets of simulations to evaluate the performance of A-TMLE in these aspects. These simulations covered a range of scenarios, including simple to complex bias functions, varying sample sizes of external data relative to the RCT, and practical violations of trial enrollment positivity. Descriptions of the five estimators we assessed are shown in Table 2.
@@
Our primary metrics for evaluating the performance of the estimators include mean-squared-error (MSE), relative MSE, coverage and average width of 95% confidence intervals. All data generating processes are detailed in Appendix B.
@@
Figure 1 presents results for scenarios with simple bias functions and varying external data sample sizes relative to the RCT. In scenarios (a) and (b), the bias functions (i.e., the conditional effect of the study indicator $S$ on the outcome $Y$) have simple main-term linear forms. Directly pooling the two data sources would yield a biased estimate of the true ATE due to the presence of the bias function. In the RCT, the probability of assignment to the treatment group is 0.67, which is common in settings where a drug has been approved and the focus is on assessing safety or secondary endpoints (also known as “asymmetric allocation”). The external dataset comprises both treatment and control arms patients. Treatment assignment in the external data is determined by baseline patient characteristics, as would be the case in routine clinical practice. To reflect different stages of a drug’s post-market lifecycle and varying availability of external real-world data, we vary the external data sample sizes from one to three times that of the RCT.
@@
Figure 1: Comparison of MSE, efficiency gain, and 95% confidence interval coverage for A-TMLE, ES-CVTMLE, PROCOVA, and TMLE estimators across increasing sample sizes of both the RCT and external data. Efficiency gain is calculated as the ratio of the MSE of the reference estimator (red dashed line, representing the RCT-only estimator) to the MSE of other estimators. The plots in the right-most column show the average proportion of cross-validation folds in which the ES-CVTMLE estimator pooled the external data, averaged across Monte Carlo runs for each sample size. Scenarios (a) and (b) represent two different data-generating-processes involving main-term parametric models for the conditional average RCT-enrollment effect.
We observe that the A-TMLE estimator consistently achieves the lowest MSE compared to alternative estimators. As the external data sample size increases from one to three times that of the RCT, the efficiency gain relative to a regular TMLE applied to the RCT data alone grows from approximately 1.1 to 2 times. All estimators show roughly 95% confidence interval coverage. The ES-CVTMLE method partitions the data into $V$ folds. For each fold, it estimates the MSE on the larger $(V-1)$-fold training set, then data-adaptively selects either the $\tilde{\Psi}$ or $\Psi$ based on the magnitude of MSE and estimates the chosen target parameter on the held-out fold. The final estimate is obtained as the average across folds. For ES-CVTMLE, we plotted the average proportion of folds that pooled the external data across Monte Carlo runs in the last column of Figure 1.
@@
In scenario (a), the bias introduced by pooling external data is relatively small compared to scenario (b). This smaller bias allows ES-CVTMLE to achieve moderate efficiency gains over regular TMLE, with the pooled parameter selected in about 30-70% of its cross-validation folds on average. This enables ES-CVTMLE to leverage external data for efficiency improvements. However, in scenario (b), the bias magnitude introduced by the external data is significantly larger, nearly eliminating ES-CVTMLE’s efficiency gain. Notice that ES-CVTMLE rejects external data in most cross-validation folds due to the large estimated bias. In contrast, A-TMLE effectively learns the simple bias function in this scenario, achieving efficiency gains while maintaining valid type I error control. Recall that in the Introduction section, we mentioned that an efficient estimator for the target parameter $\Psi$ may still offer no gain in the presence of external data, this is evident in the relative MSE plots with a regular TMLE that goes after $\Psi$ matching the MSE of a regular TMLE applied to the RCT data alone (red dashed line). Therefore, even theoretically optimal estimators may not yield efficiency gains from external data integration in finite sample.